OpenStudy (vishweshshrimali5):

tan x + cot x = 2 then find sin x .

OpenStudy (vishweshshrimali5):

\[\cfrac {1}{cot x} + cot x = 2\] \[cot^2 x - 2cot x + 1 = 0\] \[cot x = 1\] \[cosec^2 x = 1+ cot^2 x = 2\] \[cosec x = \pm \sqrt{2}\] but answer given is sin x = \(\cfrac{1}{\sqrt{2}}\) should not it be sin x = \(\cfrac{1}{\pm\sqrt{2}}\)

OpenStudy (vishweshshrimali5):

@estudier @experimentX

OpenStudy (vishweshshrimali5):

@goformit100

OpenStudy (experimentx):